Inspection costs at one of Ratulowski Corporation's factories are listed below:

Units Produced Inspection Costs

April 906 $ 16,312

May 951 $ 16,700

June 912 $ 16,365

July 896 $ 16,210

August 918 $ 16,394

September 903 $ 16,280

October 920 $ 16,432

November 860 $ 15,850

December 899 $ 16,238

Management believes that inspection cost is a mixed cost that depends on units produced.

Calculate the following using the high-low method:

1. Variable Cost per unit produced (round to 2 decimal place) $

2. Overall Fixed Cost portion of Inspection Costs (round to 2 decimal place) $

3. Cost equation for Inspection Costs (in the linear equation format) Y=

4. Estimated Inspection costs when 973 units are produced (round to nearest whole dollar) $
Answer #1

1) Variable cost per unit produced = Change in cost/Change in unit produced

                                                 = (16700-15850)/(951-866)

Variable cost per unit produced = $9.34 per unit produced

2) Overehad fixed cost portion = 16700-(951*9.34) = $7817.66

3) Cost equaiton = Y = a+bx ; Y = 7817.66+9.34X

4) Inspection cost = 7817.66+(9.34*973) = $16905